# Aligning Large Language Models with Self-generated Preference Data

### v1

@@ 5 Experiments
    Does Selfee improve the alignment of LLMs only using a small amount of human-labeled preference data? (Table 1, Figure 4)
@@
    Is Selfee generalizable across various LLMs? (Table 4)
@@
    Is Selfee also effective to improve overall LLM’s capability? (Table 6)
@@ 5.1 Experimental Setups
Models. When there are no specific mentions, our experiments were conducted using the supervised fine-tuned Mistral-7b-0.1 model [15], as the initial model $\pi_{\text{init}}$ in Section 4. Specifically, we use the open-sourced model555alignment-handbook/zephyr-7b-sft-full that follows the recipe of zephyr [40] and fine-tuned on the instructions of Ultrachat [8]. More details are presented in Appendix B.
@@
Baselines. To evaluate the effectiveness of the proposed preference judgment method (Eq. 7), we compare it with other preference judgment methods. Specifically, we consider the baselines that train the model via iterative DPO [35, 47], which iteratively generate preference data and update the model, using LLM-as-judge [3, 50] (i.e., in-context learning) or an external powerful reward model (PairRM [16]) for the preference judgment. Notably, these approaches are the same in the case of changing the judgment method and removing self-refinement in Selfee. Details are presented in Appendix B.
@@
Datasets. For the preference learning, we utilized UltraFeedback dataset [7], following the previous works [35, 34].666"argilla/ultrafeedback-binarized-preferences-cleaned" To be specific, from this dataset, we first construct the seed data, consisting of 2K samples with prompts, responses, and ground truth preference labels. Then, the remaining samples are divided into subsets of 8K, 20K, and 30K samples, leaving only the prompts. These subsets were used as the prompt sets for the iteration stages 1, 2, and 3, respectively.
Evaluations. Following the common practice, we mainly evaluate each model our evaluations using AlpacaEval2.0 [9, 10, 20]. AlpacaEval2.0 approximately evaluates human preference for instruction following. Using 805 instructions from various datasets, the evaluation is conducted by comparing the response of GPT-4 [28] and the testing model to measure win rates. To mitigate the length bias of LLM’s preference [42, 50], both vanilla and length-controlled win rates are simultaneously measured.
@@
Implementation details. After the initialization stage, we conduct three rounds of data expansion with self-generated preference data. For data expansion, we sampled 2 responses independently per each prompt with a temperature of 0.7. Then, using the SFT model as the reference model, we assign the preference label (Eq. 7). The initial DPO training to obtain $\pi_{0}$ was conducted for 3 epochs on the seed dataset. Training on each subsequent iteration was carried out for 1 epoch. For the hyper-parameter $\beta$ of DPO, we used a fixed value of $\beta=0.1$. The batch size was set to 32, and the learning rate was $5\times 10^{-7}$. We employed adamw_torch optimizer and a cosine learning rate scheduler with a warm-up phase corresponding to 10% of the total training steps. For the hyper-parameters $\alpha$ and $K\%$ for Selfee, we used fixed values of $\alpha=0.1$ and $K=10$. Additionally, a warm-up phase was included in the denoising stage, with denoising activated after 20% of the total training steps had been completed. Regarding the hyper-parameters $\lambda$ for de-coupled noise detection, we utilized the progressively reduced values of 1/2, 1/4, and 1/8 for iterations 1, 2, and 3, respectively.
@@
Table 1: Main results. Evaluation results on AlpacaEval2.0 with different variants of Mistral-7B-v0.1. Here, the human label (%) indicates the proportion of used human preference data.
Models Gold Label (%) Len-control. Win Rate Win Rate vs. GPT-4 Avg. len (# chars) Mistral-7B-v0.1 - 0.17 0.50 5692 Zephyr-7b-$\alpha$ 100 10.3 8.4 1302 Zephyr-7b-$\beta$ 100 13.2 11 1444 SFT - 7.58 4.72 901 DPO 3.3 9.03 7.68 1802 Selfee (Ours) 3.3 15.39 21.13 2884
@@
Table 2: Comparison between judgment methods. Evaluation results on AlpacaEval2.0 with iteratively trained models (from SFT model) under different preference judgment methods.
@@
Methods External Model Len-control. Win Rate Win Rate vs. GPT-4 Avg. len (# chars) Iterative DPO (PairRM) 11.87 9.46 1292 Iterative DPO (LLM-as-judge) 9.28 9.18 1711 Selfee (no refine) 14.41 19.91 3204
@@ 5.2 Main results
After completing 3 iterations of data expansion and fine-tuning, our model achieved a 21.13% win rate against GPT-4 on the AlpacaEval2.0 benchmark, as presented in Table 1. This represents a significant improvement compared to the 7.68% (7.68% $\rightarrow$ 21.13%) win rate achieved when using only 3.3% of labeled data with the standard DPO training, while the length-control win rate is also improved. (9.03% $\rightarrow$ 15.39%). Comparing our model to the Zephyr-7b series, which uses the same base model (Mistral-7B-0.1v) and is trained on the same SFT dataset and the entire UltraFeedback preference dataset, our framework achieved superior performance in both win rate (11% vs 21.13%) and length-control win rate (13.2% vs 15.39%). However, at the same time, it is noteworthy that the average length of model outputs increased from 1802 to 2884 characters through the iterations. While such length increases are commonly observed in DPO-based algorithms and likely contributed to the higher win rate, the significant improvements in length-control win rate affirm the overall enhancement in performance from Selfee.
@@
Figure 3: Improvements during iterations. Length control (LC.) win rate (%) measured by AlpacaEval2.0 is consistently improved by Selfee and it outperforms other baselines.
@@
Next, in Table 2, we present additional experimental results to validate the proposed preference judgment method. For a fairer comparison, we provide the result of the variant of Selfee (Selfee (no refine)), which removes the self-refinement procedure (in Section 4.2) and makes it as Iterative DPO with preference judgment via Eq. 7. Consequently, three experiments in Table 2 become iterative DPO variants with different preference judgment methods. Here, one can observe that our approach showed significantly better performance compared to other methods. Specifically, Selfee achieved a win rate of 19.91% against GPT-4 on AlpacaEval2.0, compared to 9.46% for the baseline with an external reward model, PairRM. In terms of length control win rate, Selfee achieved 14.41%, surpassing the reward model’s 11.84%. On the other hand, the in-context learning approach (LLM-as-judge) shows a similar win rate compared to PairRM, but falls short in length control win rate (11.87% vs 9.28%), showing the limitations of the LLM-as-judge approach. Overall, the results reveal the superiority of our direct preference judgment over other judgment methods. This superiority is consistently observed through the iterations, as shown in Figure 3.
Table 3: Ablation study. Evaluation results on AlpacaEval2.0 with iteratively trained models (from SFT) under different methodological configurations of Selfee. DE, SR, DND are abbreviations of data expansion, self-refinement, and de-coupled noise detection, respectively.
Methods DE SR DND Len-control. Win Rate Win Rate vs. GPT-4 Avg. len (# chars) SFT - - - 7.58 4.72 901 DPO - - - 9.03 7.68 1802 14.41 19.91 3204 Selfee (Ours) 14.7 19.94 3043 15.39 21.13 2884
@@ 5.3 More analyses
Ablation study. To evaluate the impact of the self-refinement components, we conducted ablation experiments by excluding both self-refinement (SR) and decoupled noise detection (DND) from the existing framework. The results are presented in Table 3. With self-refinement without decoupled noise detection (Eq. 10), we observed a slight performance improvement, with the win rate against GPT-4 marginally increasing from 19.91% to 19.94%, and the length control win rate rising from 14.41% to 14.7%. But, when the decoupled noise detection is incorporated into the self-refinement (Eq. 11), we observed significant improvements, with the win rate increasing from 19.91% to 21.13% and the length control win rate improving from 14.41% to 15.39%. Furthermore, the inclusion of the self-refinement components led to a reduction in the average response length, decreasing from 3204 to 3043 and finally to 2884 characters (i.e., improvement in length bias). Also, these results confirm that the self-refinement component is a crucial factor in enhancing performance, contributing to both higher win rates and better length control.
Compatibility with different models. Next, to verify the compatibility of our framework across different models, we conducted experiments using the Phi-2 model [21]. [^7] https://huggingface.co/microsoft/phi-2 Specifically, we conducted experiments based on the supervised fine-tuned Phi-2888lole25/phi-2-sft-ultrachat-full, which, like Mistral, has been supervised fine-tuned on the UltraChat dataset. While maintaining the core structure of our framework, we made slight adjustments to the learning rate to accommodate the different characteristics of the Phi-2 ($5\times 10^{-6}$). In addition, due to the smaller size of the Phi-2, we observed that performance improvements were not evident beyond iteration 2. Therefore, we present the results of iteration 1. As shown in Table 4, the experimental results showed that applying Selfee to Phi-2 yields a significant increase in performance. Specifically, the win rate improved from 5.67% to 9.43%, and the length control win rate increased from 7.02% to 9.1%, compared to Phi-2 trained with DPO. These results demonstrate that Selfee is effective across different models, although its impact may be limited when applied to smaller models.
@@
Figure 4: Improvements without seed data. Evaluation results on AlpacaEval2.0 with Mistral-7B-instruct-v0.1 and Selfee without seed preference dataset.
@@
Dependency on seed data. In addition, we evaluated the feasibility of using our method even without seed preference data. Namely, we want to answer whether LLM can derive explicit human preference between responses, by leveraging their intrinsic knowledge learned about humans, during the previous training, such as pre-training or supervised instruction tuning (SFT). For this experiment, we used the Mistral-7b-instruct-0.1v [15] as the initial model (i.e., $\pi_{0}$) and the Mistral-7b-0.1v-base as the reference model (i.e., $\pi_{\text{init}}$) (see the initial setup in Section 4). This setup allows us to demonstrate that our framework can function effectively even in the absence of seed preference data, when the model is sufficiently fine-tuned with iterative data expansion and learning through self-refinement. As shown in Figure 4, the win rate increased from 6.31% to 9.79%, and the length-control win rate improved from 10.14% to 11.59%, after 3 iterations. Notably, unlike experiments with seed preference data, there was no significant increase in the average length of model outputs (1077 to 1461 characters). This indicates that Selfee can effectively leverage internal information to be aligned with human preference even without seed data.
@@
Table 4: Compatibility across various LLMs. Evaluation results on AlpacaEval2.0 with different variants fine-tuned from Phi-2.
@@
Models Gold Label (%) Len-control. Win Rate Win Rate vs. GPT-4 Avg. len (# chars) Phi-2 SFT - 5.88 3.78 978 Phi-2 DPO 3.3 7.02 5.67 1342 Selfee (Ours) 3.3 9.1 9.43 1861
@@
Table 5: Additional analyses with Selfee. Evaluation results on AlpacaEval2.0 with models that fine-tuned with different judgment methods, from the resulting model of 1st iteration of Selfee.
@@
Models Len-control. Win Rate Win Rate vs. GPT-4 Avg. len (# chars) Selfee after Iteration 1 10.57 11.89 2147 Eq. 7 with initial SFT model (Ours) 15.08 19.56 2772 Eq. 7 with previous model 13.73 17.66 2757 Judgment with PairRM 13.57 13.72 1917 Judgment without reference model 12.83 12.35 1768
@@
The results are presented in Table 5. Here, the experimental results demonstrated that the method used in Selfee, where the SFT model was utilized as the reference model for preference judgment, achieved the highest performance increase. Specifically, using the model from the previous iteration as the reference model showed lower performance, with a relatively larger decrease in the length control win rate (15.08% vs 13.73%) compared to the win rate (19.56% vs 17.66%). Despite these decreases, it still outperforms using PairRM. These results may imply the importance of judging the preference through the training LLM rather than the external model, as it is less suffering from the distribution mismatch. However, without reference model (i.e., only using the likelihood of the current model), the performance increase was the lowest compared to all other cases. These findings underscore the substantial impact of the choice of proper judgment method and reference model.
Evaluation on MT-Bench. We further evaluate trained models using MT-Bench [16] to assess different aspects of LLMs. Namely, MT-Bench evaluates a chatbot’s overall abilities across multiple categories related to key LLM capabilities such as math, coding, roleplay, writing, etc. The evaluation is conducted by scoring responses to multi-turn questions using GPT-4. As shown in Table 6, Selfee achieved a score of 6.94 on the MT-Bench, outperforming other baselines and the model trained with DPO on 3.3% gold labeling data. In addition, when we refine the preference using relatively less aligned predictions (i.e., $\lambda=0$), denoted by Selfee, it even outperforms the strong baseline with the external reward model. Moreover, we observe that Selfee consistently improves the MT-bench performance without seed preference data or with smaller LLM, as presented in Table 7 in Appendix C. These results indicate that Selfee not only excels in specific tasks but also contributes to the overall performance enhancement of LLMs.
Table 6: MT-Bench. Evaluation results on MT-bench with different variants of Mistral-7B-v0.1.
Models Gold Label (%) Avg Score Mistral-7B-v0.1 - 3.25 SFT - 6.34 DPO 3.3 6.81 Iterative DPO (PairRM) 3.3 6.98 Iterative DPO (LLM-as-judge) 3.3 6.67 Selfee (Ours) 3.3 6.94 Selfee (Ours) 3.3 7.09

### v2

@@ 5 Experiments
    Does SPA improve the alignment of LLMs only using a small amount of human-labeled preference data? (Table 1, Figure 4)
@@
    Is SPA generalizable across various choices of seed data and types of LLMs? (Tables 3,4,5)
@@
    What is the effect of each component in SPA? (Tables 6,7)
@@ 5.1 Experimental setups
Models. When there are no specific mentions, our experiments were conducted using the supervised fine-tuned Mistral-7b-0.1 model (Jiang et al., 2023a), as the initial model $\pi_{\text{init}}$ in Section 4. Specifically, we use the open-sourced model666alignment-handbook/zephyr-7b-sft-full that follows the recipe of Zephyr (Tunstall et al., 2023) and fine-tuned on the instructions of Ultrachat (Ding et al., 2023). More details are in Appendix B.
@@
Baselines. To evaluate the effectiveness of the proposed preference judgment method (Eq. 7), we compare it with other preference judgment methods. Specifically, we consider the baselines that train the model via Iterative DPO (Snorkel, 2024; Xu et al., 2023), which iteratively generate preference data and update the model, using LLM-as-judge (Bai et al., 2022b; Zheng et al., 2023) (i.e., in-context learning) or an external powerful reward model (PairRM (Jiang et al., 2023b)) for the preference judgment. Notably, these approaches are the same in the case of changing the judgment method and removing self-refinement in SPA. Details are presented in Appendix B.
@@
Datasets. For the preference learning dataset, we utilized UltraFeedback (Cui et al., 2023), following the previous works (Snorkel, 2024; Rosset et al., 2024).777"argilla/ultrafeedback-binarized-preferences-cleaned" To be specific, from this dataset, we first construct the seed data, consisting of 2K samples (3.3% of 60K) with prompts, responses, and ground truth preference labels. We refer the ground-truth preference label provided by the UltraFeedback as gold label in Tables 1 and 5. Then, the remaining samples are divided into subsets of 8K, 20K, and 30K samples, leaving only the prompts. These subsets were used as the prompt sets for the iteration stages 1, 2, and 3, respectively. Only for the experiments in Table 3, the size of seed data is changed.
Evaluations. Following the common practice in LLM alignment, we mainly evaluate each model our evaluations using (1) AlpacaEval 2.0 (Dubois et al., 2023; 2024; Li et al., 2023a). AlpacaEval 2.0 approximately evaluates human preference for instruction following. Using 805 instructions from various datasets, the evaluation is conducted by comparing the response of GPT-4 (OpenAI, 2023) and the testing model to measure win rates. To mitigate the length bias of LLM’s preference (Wang et al., 2023b; Zheng et al., 2023), both original and length-controlled (LC) win rates are simultaneously measured. LC win rate is an adjusted win rate by neutralizing the effect of response length to focus on quality, using a separately trained regression model (Dubois et al., 2024). We also evaluate trained LLMs using (2) MT-Bench (Zheng et al., 2023) to assess different aspects of LLMs. Namely, MT-Bench evaluates a chatbot’s overall abilities across multiple categories related to key LLM capabilities such as math, coding, roleplay, writing, etc. The evaluation is conducted by scoring responses to multi-turn questions using GPT-4. These benchmarks also provide a thorough evaluation of LLMs’ alignment with human preferences and their overall effectiveness in practical applications.
@@
Implementation details. After the initialization stage, we conduct three rounds of data expansion with self-generated preference data. For data expansion, we sampled 2 responses independently per each prompt with a temperature of 0.7. Then, using the SFT model as the reference model, we assign the preference label (Eq. 7). The initial DPO training to obtain $\pi_{0}$ was conducted for 3 epochs on the seed dataset. Training on each subsequent iteration was carried out for 1 epoch. For the hyper-parameter $\beta$ of DPO, we used a fixed value of $\beta=0.1$. The batch size was set to 32, and the learning rate was $5\times 10^{-7}$. We employed AdamW optimizer and a cosine learning rate scheduler with a warm-up phase corresponding to 10% of the total training steps. For the hyper-parameters $\alpha$ and $K\%$ for SPA, we used fixed values of $\alpha=0.1$ and $K=10$. Additionally, a warm-up phase was included in the denoising stage, with denoising activated after 20% of the total training steps had been completed. Regarding the hyper-parameters $\lambda$ for de-coupled noise detection, we utilized the progressively reduced values of 1/2, 1/4, and 1/8 for iterations 1, 2, and 3, respectively.
@@
Table 1: Main results. Evaluation results on AlpacaEval 2.0 and MT-Bench with different variants of Mistral-7B-v0.1. The best scores are highlighted with bold.
AlpacaEval 2.0 MT-Bench
Models
Gold
Label (%)
Len-control.
Win Rate (%)
Win Rate
vs. GPT-4 (%)
Avg. Score
(0-10)
Mistral-7B-v0.1 - 0.17 0.50 3.25
Zephyr-7b-$\beta$ 100 11.75 10.03 6.87
SFT - 7.58 4.72 6.34
DPO 3.3 9.03 7.68 6.81
SPA (Ours) 3.3 15.39 21.13 6.94
@@
Table 2: Comparison with baselines for preference judgment. Evaluation results on AlpacaEval 2.0 and MT-Bench with iteratively trained models (from SFT model) under different preference judgment methods. The best scores are highlighted with bold.
@@
AlpacaEval 2.0 MT-Bench Methods External Model Len-control. Win Rate (%) Win Rate vs. GPT-4 (%) Avg. Score (0-10) Iterative DPO (PairRM) 11.87 9.46 6.98 Iterative DPO (LLM-as-judge) 9.28 9.18 6.67 SPA (Ours) 15.39 21.13 6.94
@@ 5.2 Main results
After completing 3 iterations of data expansion and fine-tuning via SPA, the trained model achieved a 21.13% win rate against GPT-4 on the AlpacaEval 2.0 benchmark, as presented in Table 1. This represents a significant improvement compared to the 7.68% (7.68% $\rightarrow$ 21.13%) win rate achieved when using only 3.3% of labeled data with the standard DPO training, while the length-control win rate is also improved. (9.03% $\rightarrow$ 15.39%). In addition, SPA achieved a score of 6.94 on the MT-Bench, clearly outperforming the model trained with DPO (6.81) on the same amount of 3.3% gold labeling data. More interestingly, our framework achieved superior performance in both win rate (10.03% vs 21.13%) and length-control win rate (11.75% vs 15.39%), compared to Zephyr-7b-$\beta$ which uses same base model (Mistral-7B-0.1v) and SFT dataset but uses significantly larger labeled preference data, i.e., 100% of UltraFeedback dataset (v.s. 3.3% for SPA). These significant improvements in both win rates clearly affirm the overall enhancement in performance from SPA.
@@
Figure 3: Improvements during iterations. Length control (LC.) win rate (%) measured by AlpacaEval 2.0 is consistently improved by SPA and it outperforms other baselines.
Next, in Table 2, we present additional experimental results to validate the proposed preference judgment method. Namely, three experiments in Table 2 can be viewed as the Iterative DPO variants with different preference judgment methods. One can observe that SPA showed significantly better performance compared to other methods. Specifically, SPA achieved a win rate of 21.13% against GPT-4 on AlpacaEval 2.0, compared to 9.46% for the baseline with an external reward model, PairRM. In terms of length control win rate, SPA achieved 15.39%, surpassing the reward model’s 11.84%. Here, we conjecture that the reason why the Iterative DPO training with the proposed direct preference judgment method (using training LLM) outperforms the case with inferred labels from the external reward model is related to the distribution shift. As the iteration is increased, the distribution of the generated data with LLM is more shifted from the distribution of the seed preference data. Then, the effectiveness of the external reward model inevitably decreases, as the reward model is fixed while the generated data is increasingly distant from its training distribution. In contrast, SPA generates the preference label using the intrinsic reward model that is continuously updated for each iteration. Therefore, it less suffers from the distribution shift during the iterations, and hence could be more effective for iterative training. Regarding this, we remark on the results in Figure 3; at iteration 1, the effectiveness of both approaches is not much different. However, the gap is significantly widened at iteration 2, and it empirically supports the above rationale.
@@
On the other hand, the in-context learning approach (LLM-as-judge) shows a similar win rate compared to PairRM, but falls short in length control win rate (11.87% vs 9.28%), showing the limitations of the LLM-as-judge approach. Overall, the results reveal the superiority of our direct preference judgment over other judgment methods. Also, this superiority is consistently observed through the iterations, as shown in Figure 3.
@@ 5.3 More analyses
In this section, we conduct additional analyses of SPA by comparing the results on AlpacaEval 2.0. More comparisons on the MT-Bench and the additional experiments are presented in the Appendix.
@@
Figure 4: Improvements without seed data. Evaluation results on AlpacaEval 2.0 with Mistral-7B-instruct-v0.1 and SPA with no seed preference data.
Table 3: Different number of seed data. Evaluation results on AlpacaEval 2.0 with Mistral-7B-v0.1 trained with DPO and SPA under the different number of seed ground-truth preference labels.
Used Ground-truth Preference Data Methods 0.8% 1.7% 3.3% 10% DPO: LC Win Rate (%) 7.85 7.68 9.03 11.37 DPO: Win Rate (%) 5.53 5.49 7.68 9.32 SPA: LC Win Rate (%) 10.36 12.36 16.23 18.52 SPA: Win Rate (%) 11.34 13.72 19.94 23.79
Generalization across different numbers of seed data. Previously, we conducted the experiment by assuming that only a limited number of human preference data is initially given, e.g., 3.3% of UltraFeedback dataset. However, the effectiveness of SPA does not depend on the size of the seed preference dataset and we validate this with the additional experiments. First, we conduct the experiments by varying the portion of the seed ground-truth preference data. Specifically, to use the fixed input prompt datasets for each iteration, we consider the following portions for the experiments: [0.8%, 1.7%, 10%]. Table 3 shows the results on AlpacaEval 2.0 with Mistral-7B-v0.1 after 2 iterations of training with SPA, including the original experiments with 3.3% seed preference data. Here, one can observe that the alignment performance under DPO and SPA is improved with the increased seed data, and SPA consistently outperforms DPO which demonstrates the robustness of SPA regarding the size of seed preference data.
@@
We further evaluated the feasibility of using SPA even without seed preference data. Namely, we want to answer whether LLM can derive explicit human preference between responses, by leveraging their intrinsic knowledge learned about humans, during the previous training, such as pre-training or supervised instruction tuning (SFT). For this experiment, we used the Mistral-7b-instruct-0.1v (Jiang et al., 2023a) as the initial model (i.e., $\pi_{0}$) and the Mistral-7b-0.1v-base as the reference model (i.e., $\pi_{\text{init}}$) (see the initial setup in Section 4). This setup allows us to demonstrate that our framework can function effectively even in the absence of seed preference data, when the model is sufficiently fine-tuned with iterative data expansion and learning through self-refinement. As shown in Figure 4, the win rate increased from 6.31% to 9.79%, and the length-control win rate improved from 10.14% to 11.59%. This result indicates that SPA can leverage the internal information of LLMs to be aligned with human preference even without seed data.
Table 4: Different initial seeds. Evaluation results on AlpacaEval 2.0 with different variants of Mistral-7B-v0.1 under the different sampling of the initial seed preference data.
Methods 1st Seed Data 2nd Seed Data 3rd Seed Data Average Variance DPO: LC Win Rate (%) 9.03 8.74 9.54 9.10 0.16 DPO: Win Rate (%) 7.68 7.17 7.59 7.48 0.07 SPA (Ours): LC Win Rate (%) 16.23 13.77 16.38 15.46 2.10 SPA (Ours): Win Rate (%) 19.94 20.06 19.74 19.91 0.03
Variance with different initial seed dataset. In addition, we conduct experiments to check the sensitivity of SPA with the initial seed preference dataset by varying them with different random sampling. The results after 2 iterations of training with SPA are presented in Table 4. Here, one can observe that the proposed SPA consistently improves the alignment performance regardless of the given seed data, and the variance between them is not significant, especially in the case of a normal win rate. While ours exhibits a relatively high variance for length-controlled (LC) win rate, its lowest confidence interval value (13.36 %) is certainly higher than the value of the strongest baseline (11.98 %) which confirms the effectiveness of our method.
Compatibility with different models. Next, to verify the compatibility of our framework across various LLMs, we conducted experiments using three different LLMs: Phi-2-2.7B (Li et al., 2023b), LLaMA3-8B (Dubey et al., 2024), and Phi-3-14B. Specifically, we conducted experiments based on their supervised fine-tuned versions; for Phi-2, we used the model that has been fine-tuned on the UltraChat dataset like Mistral.888lole25/phi-2-sft-ultrachat-full For LLaMA-3999https://huggingface.co/meta-llama/Meta-Llama-3-8B-Instruct and Phi-3101010https://huggingface.co/microsoft/Phi-3-medium-4k-instruct, we used the generally fine-tuned models as there are no models that have been fine-tuned on the UltraChat dataset. Here, most of the experimental setups for these experiments are maintained, and the slightly adjusted setups are detailed in Appendix B.3. As shown in Table 5, the experimental results showed that applying SPA to various LLMs yields consistent improvements in the performance. For example, the win rate improved from 5.67% to 9.43%, and the length control win rate increased from 7.02% to 9.1%, in the case of Phi-2 after being trained with SPA compared to DPO. These results demonstrate that the effectiveness of SPA is not limited to the specific LLMs and is generalized across various LLMs.
@@
Table 5: Compatibility across various LLMs. Evaluation results on AlpacaEval 2.0 with different training methods (SFT, DPO, and SPA) across various types of LLMs (Phi-2-2.7B, LLaMA-3-8B, and Phi-3-14B). The best scores are highlighted with bold.
@@
Phi-2-2.7B LLaMA-3-8B-Instruct Phi-3-14B-Instruct Methods Gold Label (%) Len-control. Win Rate (%) Win Rate vs. GPT-4 (%) Len-control. Win Rate (%) Win Rate vs. GPT-4 (%) Len-control. Win Rate (%) Win Rate vs. GPT-4 (%) SFT - 5.88 3.78 21.40 21.71 26.51 21.41 DPO 3.3 7.02 5.67 24.17 25.39 27.70 22.12 SPA (Ours) 3.3 9.10 9.43 25.03 34.84 28.77 24.14
Table 6: Ablation study. Evaluation results on AlpacaEval 2.0 with iteratively trained models (from SFT) under different methodological configurations of SPA. DE, SR, DND are abbreviations of data expansion, self-refinement, and de-coupled noise detection, respectively. The best scores are highlighted with bold.
AlpacaEval 2.0 Methods DE SR DND Len-control. Win Rate (%) Win Rate vs. GPT-4 (%) SFT - - - 7.58 4.72 DPO - - - 9.03 7.68 14.41 19.91 SPA (Ours) 14.7 19.94 15.39 21.13
@@
Table 7: Additional analyses. Evaluation results on AlpacaEval 2.0 with models that fine-tuned with different judgment methods, from the resulting model of 1st iteration of SPA.
@@
AlpacaEval 2.0 Models Len-control. Win Rate (%) Win Rate vs. GPT-4 (%) SPA after iteration 1 10.57 11.89 Eq. 7 with initial SFT model (Ours) 15.08 19.56 Eq. 7 with previous model 13.73 17.66 Judgment with PairRM 13.57 13.72 Judgment without reference model 12.83 12.35
Ablation study. To evaluate the impact of the self-refinement components, we conducted ablation experiments by excluding both self-refinement (SR) and decoupled noise detection (DND) from the existing framework. The results are presented in Table 6. With self-refinement without decoupled noise detection (Eq. 10), we observed a slight performance improvement, with the win rate against GPT-4 marginally increasing from 19.91% to 19.94%, and the length control win rate rising from 14.41% to 14.7%. But, when the decoupled noise detection is incorporated into the self-refinement (Eq. 11), we observed significant improvements, with the win rate increasing from 19.91% to 21.13% and the length control win rate improving from 14.41% to 15.39%. Also, these results confirm that the self-refinement component is a crucial factor in enhancing performance, contributing to both higher win rates and better length control.
@@
The results are presented in Table 7. Here, the experimental results demonstrated that the method used in SPA, where the SFT model was utilized as the reference model for preference judgment, achieved the highest performance increase. Specifically, using the model from the previous iteration as the reference model showed lower performance, with a relatively larger decrease in the length control win rate (15.08% vs 13.73%) compared to the win rate (19.56% vs 17.66%). Despite these decreases, it still outperforms using PairRM. These results may imply the importance of judging the preference through the training LLM rather than the external model, as it is less suffering from the distribution mismatch. However, without reference model (i.e., only using the likelihood of the current model), the performance increase was the lowest compared to all other cases. These findings underscore the substantial impact of the choice of proper judgment method and reference model.